OpenStudy (anonymous):

Please help me find the Integral of (x+5) / (sqrt[9-(x-3)^2])

OpenStudy (anonymous):

\[\int\limits_{}^{}(x+5)/\sqrt{9-(x-3)^2}\]

OpenStudy (anonymous):

trying to recall the arctrigs ...

OpenStudy (anonymous):

We learned 3 arcsin,arctan, and arcsec I would say that this is arcsin

OpenStudy (anonymous):

y = sin'x siny = x cosy y' = 1 y' = 1/cos(sin'x) y' = 1/sqrt(1-sin^2(sin'x)) im thinking its arcsin as well

OpenStudy (anonymous):

So I would also say that u=x-3

OpenStudy (anonymous):

let sin(u) = x-3 du = dx

OpenStudy (anonymous):

du = 1 a=3

OpenStudy (anonymous):

... we want a 9sin^2, so let 3sin(u) = x-3

OpenStudy (anonymous):

Why 9sin^2?

OpenStudy (anonymous):

to factor out the sqrt9 to get sqrt[1-trig^2]

OpenStudy (anonymous):

Is that a rule because in my mind sqrt9 = 3

OpenStudy (anonymous):

sqrt[9-9a] = sqrt[ 9(1-a)] ] = 3 sqrt(1-a) the 1-a is the format we want to work this into

OpenStudy (anonymous):

Oh because that is in the original formula okay. Oh wait did you multiply out (x-3)^2?

OpenStudy (anonymous):

no, we want the trig squared, so its the innards that we are subbing 3sin(u) = x-3; x = 3sin(u)-3 3cos(u) du = dx \[\frac{x+5}{\sqrt{9-(x-3)^2}}dx\] \[\frac{3sin(u)-3+5}{\sqrt{9-(3sin(u))^2}}~3cos(u) du\]

OpenStudy (anonymous):

Oh you solved for x and multiplied by u' Wouldn't x=3sin(u)+3 because you're taking -3 to the other side of the equation?

OpenStudy (anonymous):

lol, yes it would :)

OpenStudy (anonymous):

Oh okay. So . . . is that it? Or do we have to get it into the arcsin formula now?

OpenStudy (anonymous):

we can simplify at the moment, that bottom goes 3 sqrt(1-sin^2) = 3cos, so we are left with: 3sin(u)+8 du as our integration

OpenStudy (anonymous):

If the integral = arcsin(u/a) + c Then is the answer arcsin(x-3)/3 + C ?

OpenStudy (anonymous):

the usub led us in a different direction i believe, or at least simplified it tremendously we can use the wolf to dbl chk

OpenStudy (anonymous):

Oh lol yeah I forgot about that website. Let me try it

OpenStudy (anonymous):

integrates to: 3cos(u) + 8u, along the interval

OpenStudy (anonymous):

or +C for the indefinite

OpenStudy (anonymous):

so, what is u? 3sin(u) = x-3 sin(u) = x/3 - 1 u = arcsin (x/3 - 1 )

OpenStudy (anonymous):

Okay thanks. I"m sorry I have to go but you've been a great help :)

OpenStudy (anonymous):

youre welcome, ill have to dbl chk my thoughts to be sure :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=d%2Fdx+%28-3cos%28arcsin+%28x%2F3+-+1+%29%29%2B8arcsin+%28x%2F3+-+1+%29%29 forgot the negative .. -3cos(u) + 8u the wolf says it derives to a sqrt(-(x-6)x) 9 - (x^2 -6x + 9) = -(x-6)x